Plot the following points:

P1 (-3, 135°)

P2 (2, )

P3 (4, 405°)

Graph

Sketch the graph of r = 3 − 2cosθ.



1
Expert's answer
2021-12-12T18:15:58-0500

1.



"P_1 (-3, 135\u00b0)"

Or



"P_1 (-3\cos(135\degree), -3\sin(135\degree))"






"P_2 (2, -3.14\/3)"

Or



"P_2 (2\cos(-3.14/3), 3\sin(-3.14/3))"






"P_3 (4, 405\degree)"

Or



"P_3 (4\cos(405\degree), 4\sin(405\degree))""r = 3 − 2\cosθ""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 \theta (rad) & 0 & \pi/6 & \pi/4 & \pi/3 & \pi/2 \\ \hline
 r & 1 & 3-\sqrt{3} & 3-\sqrt{2} & 2 & 3 \\
\end{array}""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \theta (rad) & 2\pi/3 & 3\pi/4 & 5\pi/6 & \pi \\ \hline
 r & 4 & 3+\sqrt{2} & 3+\sqrt{3} & 5 \\
\end{array}""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \theta (rad) & 7\pi/6 & 5\pi/4 & 7\pi/3 & 3\pi/2 \\ \hline
 r & 3+\sqrt{3} & 3+\sqrt{2} & 4 & 3 \\
\end{array}""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \theta (rad) & 5\pi/3 & 7\pi/4 & 11\pi/6 & 2\pi \\ \hline
 r & 2 & 3-\sqrt{2} & 3-\sqrt{3} & 1 \\
\end{array}"